OpenStudy (anonymous):

What is the derivative and the second derivatie of x^2lnx? and what are the critical points of f' and f''?

OpenStudy (anonymous):

I know to use product rule. I just need to confirm my answer.

OpenStudy (anonymous):

I'm looking at it now, but it's not the product rule, as you don't have two distinct expressions MULTIPLED together, one expression (x), is raised to the POWER of another expression (2lnx).

OpenStudy (mimi_x3):

\[x^2lnx\] \[\large 2xlnx + \frac{x^{x}}{x} = 2xln +x\] I think that its the product rule..

OpenStudy (anonymous):

Oh I"m sorry, I should have noted it better. It's (x^2)(lnx). My bad

OpenStudy (anonymous):

How would I find critical points? I know you have to make y=0

OpenStudy (anonymous):

and what is the second derivative of that?

OpenStudy (anonymous):

ok, yes, I misread your question. y = (x^2).(ln x) y' = (x^2).(1/x) + (2x).(ln x) = x + 2x.lnx y'' = 1 + 2x.(1/x) + 2lnx = 3 + 2lnx To find critical points: y = 0 gives x-intercept y' = 0 gives stationary points (turning points & horizontal inflextions) use the 2nd derivative to determine the nature of the above y''=0 gives points of inflexion

OpenStudy (anonymous):

So the only critical point for f'(x) is 0?

OpenStudy (anonymous):

f'(x) = x+2xlnx = x(1+2lnx) = 0 when x = 0, and when 1+2lnx = 0 1+2lnx = 0 2lnx = -1 lnx = -1/2 x = e^(-1/2) be thorough!

OpenStudy (anonymous):

So the two critical points are 0 and e^0.5?

OpenStudy (anonymous):

0 and e to the power MINUS 0.5

OpenStudy (anonymous):

Sorry I forgot that minus :P so for the second derivative, the critical point would be e^-3/2?

OpenStudy (anonymous):

what you refer to as "critical points" might be different to me (I'm in Australia), & we refer to intercepts, stationary points, & inflexions - all of which, personally, I would regard as "critical" to getting a good sketch. 2nd derivative: yes, you're right there.

OpenStudy (anonymous):

Critical Point: A critical point on a graph occurs at x=c if f(c) is defined, and if and only if f'(c) either is zero or is undefined

OpenStudy (anonymous):

so for the second derivative, we would set f''(x) equal to zero so 0=3+2lnx and I got x to be e^-3/2. is that right?

OpenStudy (anonymous):

and is that the ONLY critical point?